How long will it take to double my investment (which they don't tell me) with compound annual investment, compound monthly and compound continuously. please explain in detail
Answer by Alan3354(69443)

How long will it take to double my investment (which they don't tell me) with compound annual investment
FV = PV*(1 + r)^n r = interest rate per period, n = # of periods
FV = 2*PV
------
(1 + r)^n = 2
n*log(1 + r) = log(2)
n = log(2)/log(1 + r)
The value of n depends on the interest rate.
The amounts are not relevant, whatever you start with is doubled.
---------------
compound monthly:
Same as above, but the monthly interest rate = annual rate/12, and n is the # of months.
----------------
and compound continuously.
FV = PV*e^rt, r = rate per period, t = periods
e^rt = 2
rt = ln(2)
t = ln(2)/r
-----
You have to use ln (natural logs) for this one. The 1st 2 can be log or ln.
--------------
please explain in detail
FV = Future Value
PV = Present Value
-------------